OpenStudy (anonymous):

Which shows 632 - 372 being evaluated using the difference of squares method?

OpenStudy (anonymous):

632 - 372 = (3969 + 1369)(3969 - 1369) = 13,878,800 632 - 372 = 3969 - 1369 = 2,600 632 - 372 = (63 - 37)2 = 262 = 676 632 - 372 = (63 + 37)(63 - 37) = (100)(26) = 2,600

OpenStudy (anonymous):

@mipie13 can you help me?